40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 18 12:04:39 2008, Alberto, Update, Electronics, RF Monitor (StocMon) DSC_0443.JPG
    Reply  Fri Feb 22 08:29:07 2008, Alberto, Update, Electronics, RF Monitor (StocMon) 
       Reply  Fri Feb 22 11:13:15 2008, rob, Update, Electronics, RF Monitor (StocMon) 
    Reply  Fri Feb 22 08:33:18 2008, Alberto, Update, Electronics, RF Monitor (StocMon) 
       Reply  Thu Feb 28 19:49:21 2008, rob, Update, Electronics, RF Monitor (StocMon) 
Message ID: 331     Entry time: Fri Feb 22 08:29:07 2008     In reply to: 321     Reply to this: 334
Author: Alberto 
Type: Update 
Category: Electronics 
Subject: RF Monitor (StocMon) 

Quote:
I put the amplifiers next to the monitor on the PSL table, layed the power and the RF SMA cables out to the rack. I'm powering the box and the amplifiers with the power supply, waiting for someone to show me tomorrow how to connect it to the Sorensen (Steve, Ben?).

I'm ready to hook up the channels into EPICS.


Me and Ben Abbot were plugging the cables that power that RF Monitor box into the PSL rack when inadvertently we made some arcs spark between the pins on the back of one of the ADC. Somehow that made the laser shut down although the MOPA stayed on. We also notice some smell of burn.

Later on, after several failed attempts, Rob, Ben and Steve could restart the laser. It took some times because the written procedure to start the chiller is not very precise.
ELOG V3.1.3-